Sorry if this isn't the right topic.... this is from an accuplacer practice test. I know the answer is 11z-20 over z(z-5) but don't understand how to work it. Here's the problem: 4/z plus 7/z-5
Answer by london maths tutor(243)

4/z plus 7/z-5
you need to obtain common denominator before adding the numerator.
the denominator = z(z-5)
but remember to multiply the same factor to the numerator too i.e
4/z = [4*(z-5)]/[z*(z-5)]
7/z-5 = [7*z /(z*(z-5))]
So the sum becomes:
[4*(z-5) + 7*z] / [z(z-5)]
[4z - 20 + 7z] / [z(z-5)]
(11z-20)/z(z-5) (hence shown)